Freshy Milk Company produces dairy equipment. Most of its jobs have a number of units per job. The company has two different departments through which all jobs pass. Overhead is applied using a plant wide rate of $15 per direct labor hour. Direct labor wages average $6 an hour. Data for Job #3 for the year is: Direct materials $30,000 Direct labor costs Dept. A $50,000 Dept. B 10,000 Machine hours used Dept. A 300 hrs. Dept. B 3,100 hrs. Units produced 34,000 units Compute the cost per unit. (Note: Round the unit cost to two decimal places.)

a.$3.50

b.$10.05

c.$7.06

d.$6.32

e.None of these choices are correct.
Answer #1

Answer : C) $ 7.06

Total Cost of Job #3 = Direct material+ Direct Labor + Manufacturing overheads

= 30,000+50,000+10,000+150,000 = 230,000

Units cost = 240,000 /34,000 = $7.06 (Answer)

Manufacturing overhead = Plant overehead rate *Total direct labour hors

= $15 *(60,000/6) = 150,000